2. In a combustion reaction, butane (C4H10) is burned in a lighter at a rate of 1.50 mol/min.

a) What is the rate at which O2(g) is consumed and CO2(g) is produced?

b) How long does 15.0g of butane take to burn?
1
Expert's answer
2020-11-18T13:42:44-0500

4H10 + 13O2 → 8CO2 + 10H2O

a) Rate(O2) "= \frac{13}{2}Rate(С_4H_{10}) = \frac{13}{2}1.50 = 9.75 \;mol/min"

Rate(CO2) "= \frac{8}{2}Rate(С_4H_{10}) = \frac{8}{2}1.50 = 6.0 \;mol/min"

b) m = 15.0 g

M(butane) = 58.12 g/mol

n(butane) "= \frac{15.0}{58.12} = 0.258 \;mol"

Time "= \frac{n}{Rate} = \frac{0.258}{1.50} = 0.172\;min = 10.3 \;sec"